OpenStudy (anonymous):

For which system of equations would elimination be the simplest technique to choose? A. y = 2x and x + 3y = 5 B. 4x + 3y = 6 and 3y = 5 C. x + y = 7 and x - y = 12 D. 3x - 2y = 8 and 4x - 5y = 20

OpenStudy (aroub):

C. x + y = 7 and x - y = 12

Parth (parthkohli):

The third one because you can easily subtract both equations.

Parth (parthkohli):

Or I'd say add.

OpenStudy (anonymous):

Aren't you suppose to explain rather than give answers?

OpenStudy (anonymous):

Yes guys, please explain.

Parth (parthkohli):

I did, didn't I?

OpenStudy (aroub):

Yeah, sorry. Parth did though.

Parth (parthkohli):

In the third one, you see, that you can eliminate y by adding equation {1} and {2}. Then you can solve for x.

OpenStudy (anonymous):

I meant that for Aroub, sorry Parth.

Parth (parthkohli):

When solved for x, then you can substitute it back to the equation, and then solve for y.

Parth (parthkohli):

You should use the method of elimination when you have the same variable with the same numerical coefficient.

OpenStudy (anonymous):

C. x + y = 7 and x - y = 12